OpenStudy (anonymous):

The radius of a circular puddle is growing at a rate of 15 cm/s. (a) How fast is its area growing at the instant when the radius is 30 cm?

OpenStudy (badhi):

The area of a circle when the radius is r, $$A=\pi r^2$$ differentiating wrt t- time, $$\frac{dA}{dt}=\pi \frac{dr^2}{dt}=\pi 2r\frac{dr}{dt}$$ substitute the values

OpenStudy (anonymous):

it says thats not right :(

OpenStudy (badhi):

can you post your steps that you've tried?